# Non-Rigid Designators in Modal and Temporal Free Description Logics

### v1

@@ 4 Non-Rigid Designators and Counting
Note that our language contains existential quantification, which in (?) is introduced as an abbreviation for $C\sqcup\exists^{\neq}u.C$. $\mathcal{ML}_{\text{{Diff}}}^{n}$ can be regarded as a basic first-order modal logic with counting because the counting quantifier $\exists^{=1}u.C$ with
@@
is logically equivalent to the $\mathcal{ML}_{\text{{Diff}}}^{n}$-concept $\exists u.(C\sqcap\neg\exists^{\neq}u.C)$. Conversely, $\exists^{\neq}u.C$ is logically equivalent to $\exists u.C\sqcap(C\Rightarrow\neg\exists^{=1}u.C)$. So, one could replace $\exists^{\neq}$ by $\exists^{=1}$ in the definition $\mathcal{ML}_{\text{{Diff}}}^{n}$ and obtain a logic with exactly the same expressive power.
@@
(1) $\mathcal{C}$-satisfiability of $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concepts (under global ontology) can be reduced in double exponential time to $\mathcal{C}$-satisfiability of $\mathcal{ML}_{\text{{Diff}}}^{n}$-concepts (under global ontology, respectively), both with constant and expanding domains.
@@
The proof of (1) and proofs in Sections 5 and 6 use types ${\boldsymbol{t}}$, sets of (negated) subconcepts of the input concept (and ontology) such that, for any subconcept $C$, either $C\in{\boldsymbol{t}}$ or $\neg C\in{\boldsymbol{t}}$. A quasistate $T$ is a set of types and stands for a world satisfying exactly the types in $T$ (?). Given a quasistate $T$, one can check in polytime whether its non-modal abstraction, obtained by replacing all concepts starting with a modal operator by a concept name, is satisfiable in a DL interpretation. On the other hand, assuming that $\mathcal{C}$-satisfiability of $\mathcal{ML}_{\text{{Diff}}}^{n}$-concepts is decidable, one can check whether its non-DL abstraction, in which existential restrictions and nominals are replaced by concept names, is satisfiable in a $\mathcal{ML}_{\text{{Diff}}}^{n}$ model in which the concept names replacing nominals are interpreted as singletons. As observed above, to enforce interpretation of a concept name $A$ as a singleton in all worlds one can use the $\mathcal{ML}_{\text{{Diff}}}^{n}$ CI $\top\sqsubseteq\exists^{=1}u.A$. For concept satisfiability (without ontology) one makes the same statement for all worlds reachable along relevant paths. The reduction is now implemented using both abstractions. It requires double exponential time because the number of quasistates is double exponential in the size of the input. Item (2) is proved by lifting to the modal description logic setting the observation of ? (?) that the difference modality and nominals are mutually interpretable by each other; see the equivalences that show the same expressive power of $\exists^{\neq}$ and $\exists^{=1}$.
Using Theorem 10, one can transfer a large number of (un)decidability results and lower complexity bounds from first-order modal logics with ‘elsewhere’ to modal DLs with non-rigid designators. Conversely, the results proved below entail new results for first-order modal logics with ‘elsewhere’ and/or counting.
@@ 5 Reasoning in Modal Free Description Logics
The aim of this section is to show the results presented in Table 1 and also discuss a few related results. For some basic frame classes, it will be convenient to use standard modal logic notation when discussing the satisfiability problem. So, given a propositional modal logic $L$ with $n$ operators, the class $\mathcal{C}_{L}$ of frames comprising all frames validating $L$, and a DL fragment $\mathcal{DL}$, then $L_{\mathcal{DL}}$ concept satisfiability (under global ontology) is the problem to decide $\mathcal{C}_{L}$-satisfiability of $\mathcal{ML}^{n}_{\mathcal{DL}}$-concepts (under global ontology, respectively). We focus on the following logics $L$ characterised by classes $\mathcal{C}_{L}$ of frames, with $n\geq 1$:
@@
$\mathcal{C}_{L}$ is as for $\mathbf{K}^{\ast n}$ and, in addition, $W$ is finite and $R$ is irreflexive (in other words, there is no chain $w_{0}R_{i_{1}}w_{1}\cdots R_{i_{n}}w_{n}$ with $w_{0}=w_{n}$).
@@
For $L\in\{\mathbf{K}^{n},\mathbf{S5}^{n}\}$ with $n\geq 1$, $L_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ concept satisfiability is in NExpTime with both expanding and constant domains.
@@
$\vec{w}=w_{0}i_{0}w_{1}\cdots i_{m-1}w_{m}$
where $1\leq i_{j}\leq n$, $i_{j}\not=i_{j+1}$, and each $R_{i}$ is the smallest equivalence relation containing all pairs of the form $(\vec{w},\vec{w}iw)\in W\times W$. One can assume that $m$ is smaller than the modal depth of $C$. Next, one can substitute the first-order domain by quasistates (as introduced above) and work with quasimodels $\mathfrak{Q}=(\mathfrak{F},\boldsymbol{q},\mathfrak{R})$, in which $\boldsymbol{q}$ associates with any world a quasistate, and a set of runs $\mathfrak{R}$ represents first-order domain elements as functions mapping every world to a type in $\boldsymbol{q}(w)$. If ${\boldsymbol{t}}\in\boldsymbol{q}(w)$ contains a nominal, there is only one $r\in\mathfrak{R}$ with $r(w)={\boldsymbol{t}}$. Now, one can apply selective filtration and some surgery to such a quasimodel to obtain a quasimodel of the above form with at most exponential outdegree (and so of at most exponential size), from which one can then reconstruct a model of exponential size.
@@
[…]
@@
Concept satisfiability for $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{LTL}_{\smash{\mathcal{ALCO}_{\!u}}}$ are polytime-reducible to concept satisfiability for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{K}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$, respectively, with and without ontology and under constant and expanding domains.
The proof of this reduction from logics of linear frames with transitive closure to logics of branching frames with transitive closure is not trivial but can be done by adapting the reduction given in the proof of Theorem 6.24 in (?) for product modal logics. Finally, the undecidability of concept satisfiability for $\mathbf{K}^{n}$ and $\mathbf{S5}^{n}$ under ontologies with constant domain can be proved in the same way as the undecidability of $\mathcal{ML}_{\text{{Diff}}}^{n}$ extended with the universal modality over arbitrary frames shown in (?) using the reduction of Theorem 10.
For many important modal logics the decidability status of modal DLs with non-rigid designators remains open. Most prominently, for the modal logics of (reflexive) transitive frames ${\bf K4}$ (and ${\bf S4}$, respectively), decidability of concept satisfiability with or without ontologies and with expanding or constant domains is open. As a “finitary” approximation and a first step to understand ${\bf K4}$ and ${\bf S4}$, we prove, as a specialisation of the proof of Theorem 12, decidability of concept satisfiability for the Gödel-Löb logic $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ ($\mathbf{GL}$ is the logic of all transitive and Noetherian111$(W,R)$ is Noetherian if there is no infinite chain $w_{0}Rw_{1}R\cdots$ with $w_{i}\not=w_{i+1}$. frames (?)) and Grzegorczyk’s logic $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ ($\mathbf{Grz}$ is the logic of all reflexive and transitive Noetherian frames (?)) in expanding domain models with and without ontologies. Alternatively, the decidability of concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ and $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ can be proved by a double exponential time reduction (similar to Theorem 10) to satisfiability in expanding domain products of transitive Noetherian frames, which is known to be decidable (?).
@@ 6 Reasoning in Temporal Free Description Logics
A flow of time $\mathfrak{F}$ is a pair $(T,<)$, where $T$ is either the set $\mathbb{N}$ of non-negative integers or a subset of $\mathbb{N}$ of the form $[0,n]$, for $n\in\mathbb{N}$, and $<$ is the strict linear order on $T$. Elements of $T$ are called instants (rather than worlds). A flow of time $(T,<)$ naturally gives rise to a frame $(T,<,\textit{succ})$ for $\smash{\mathcal{TL}_{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$, where succ is the successor relation: $\textit{succ}=\{(t,t+1)\mid t,t+1\in T\}$. So, we will often say that an interpretation $\mathfrak{M}$ is based on a flow $\mathfrak{F}$ if its frame is induced by $\mathfrak{F}$. If $\mathfrak{M}$ is based on $(\mathbb{N},<)$, then we call it an $\smash{\mathbf{LTL}_{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ interpretation and often denote (with an abuse of notation) by $\mathfrak{M}=(\Delta,(\mathcal{I}_{t})_{t\in\mathbb{N}})$; whereas, if it is based on $([0,n],<)$, for some $n\in\mathbb{N}$, it is called an $\smash{\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\!\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ interpretation and often denoted by $\mathfrak{M}=(\Delta,(\mathcal{I}_{t})_{t\in[0,n]})$.
@@
Note that $\Diamond$ is interpreted by $<$ and thus does not include the current instant.
@@
$\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{LTL}_{\smash{\mathcal{ALCO}_{\!u}}}$ $\Sigma_{1}^{1}$-complete undecidable $\Sigma_{1}^{1}$-complete undecidable
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}_{\!u}}}$ undecidable
@@
$\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}}}$ and $\mathbf{LTL}_{\smash{\mathcal{ALCO}}}$ $\Sigma_{1}^{1}$-complete undecidable
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ALCO}}}$ and $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}}}$ undecidable
@@
$\mathbf{LTL}^{\textstyle\circ}_{\smash{\mathcal{ALCO}_{\!u}}}$ / $\mathbf{LTL}^{\textstyle\circ}_{\smash{\mathcal{ALCO}}}$ Exp-complete / in Exp Exp-complete / in Exp undecidable ?
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\textstyle\circ}_{\smash{\mathcal{ALCO}_{\!u}}}$ / $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\textstyle\circ}_{\smash{\mathcal{ALCO}}}$ Exp-complete / in Exp Exp-complete / in Exp undecidable decidable
@@
We begin the study of the satisfiability problems for temporal DLs based on $\smash{\mathcal{ALCO}}$ by showing that concept satisfiability in constant domains is undecidable or even $\Sigma_{1}^{1}$-complete over the infinite flow of time $(\mathbb{N},<)$. This is very different from the classical case with RDA, which was shown to be decidable in the absence of definite descriptions (?, Theorem 14.12).
@@
The lower bounds follow, using Theorem 10 (2), from the respective undecidability results for the first-order one-variable temporal logic with counting (?). They could also be proven more directly by encoding the $\Sigma_{1}^{1}$-complete recurrence and undecidable reachability problems for the Minsky counter machines (??). Intuitively, the value of each counter of the Minsky machine can be represented as the cardinality of a certain concept. Then, non-rigid nominals (or indeed the counting quantifier) can be used to ensure that the value of the counter is incremented/decremented (depending the command) by the transition: for instance, CI of the form
@@
(1) With constant domains, $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ELO}}}$ CI entailment is $\Pi_{1}^{1}$-complete and $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ELO}}}$ CI entailment is undecidable.
(2) With expanding domains, $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ELO}}}$ CI entailment is undecidable.
(3) With expanding domains, $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ELO}}}$ CI entailment is decidable but Ackermann-hard.

### v2

@@ 4 Non-Rigid Designators and Counting
Note that our language contains existential quantification, which in (?) is introduced as an abbreviation for $C\sqcup\exists^{\neq}u.C$. In fact, $\mathcal{ML}_{\text{{Diff}}}^{n}$ can be regarded as a basic first-order modal logic with counting because the counting quantifier $\exists^{=1}u.C$ (‘there is exactly one $C$’) with
@@
is clearly logically equivalent to the $\mathcal{ML}_{\text{{Diff}}}^{n}$-concept $\exists u.(C\sqcap\neg\exists^{\neq}u.C)$. Conversely, $\exists^{\neq}u.C$ is logically equivalent to $\exists u.C\sqcap(C\Rightarrow\neg\exists^{=1}u.C)$. So, one could replace $\exists^{\neq}$ by $\exists^{=1}$ in the definition $\mathcal{ML}_{\text{{Diff}}}^{n}$ and obtain a logic with exactly the same expressive power.
@@
(1) $\mathcal{C}$-satisfiability of $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concepts (under global ontology) can be reduced in double exponential time to $\mathcal{C}$-satisfiability of $\mathcal{ML}_{\text{{Diff}}}^{n}$-concepts (under global ontology, respectively), with both constant and expanding domains.
@@
We first give the main ingredients for the proof of Item (1) with global ontology $\mathcal{O}$. If $\mathcal{O}$ contains no roles apart from $u$, then we introduce a surrogate concept name $\{a\}^{\sharp}$ for each individual name $a$ in $\mathcal{O}$ and denote by $\mathcal{O}^{\sharp}$ the result of replacing each $\{a\}$ with $\{a\}^{\sharp}$ and then extending the ontology with CIs of the form $\top\sqsubseteq\exists^{=1}u.\{a\}^{\sharp}$. The resulting $\mathcal{ML}_{\text{{Diff}}}^{n}$ ontology $\mathcal{O}^{\sharp}$ is clearly a model conservative extension of $\mathcal{O}$, which completes the reduction. If, however, $\mathcal{O}$ contains role names, then we apply the quasimodel technique (?) to deal with binary relations (roles), whose interpretations in different worlds are independent from each other. This technique is also used in Sections 5 and 6. In quasimodels, each DL interpretation is represented as a quasistate, which is a non-empty set of types, maximal consistent sets of subconcepts of $\mathcal{O}$ and their negations: each type ${\boldsymbol{t}}$ represents all domain elements satisfying the concepts in ${\boldsymbol{t}}$. In this proof, we can characterise possible quasistates for $\mathcal{O}$ using concepts: the description $\Xi_{\boldsymbol{T}}$ of a quasistate $\boldsymbol{T}$ is
$\forall u.\bigsqcup_{{\boldsymbol{t}}\in\boldsymbol{T}}{\boldsymbol{t}}\ \ \sqcap\ \ \bigsqcap_{{\boldsymbol{t}}\in\boldsymbol{T}}\exists u.{\boldsymbol{t}},$
where ${\boldsymbol{t}}$ is the conjunction ($\sqcap$) of all concepts $C\in{\boldsymbol{t}}$. The set $\mathcal{S}_{\mathcal{O}}$ of quasistates that can possibly occur in interpretations satisfying $\mathcal{O}$ can be obtained by checking whether the modal abstraction $\Xi_{\boldsymbol{T}}^{\ast}$ of its description $\Xi_{\boldsymbol{T}}$ is satisfiable, where $\cdot^{\ast}$ is the result of replacing subconcepts starting with a modal operator by fresh concept names. Since satisfiability of $\smash{\mathcal{ALCO}_{\!u}}$-concepts under $\smash{\mathcal{ALCO}_{\!u}}$ ontologies is in ExpTime, the set $\mathcal{S}_{\mathcal{O}}$ can be computed in double exponential time. In order to ensure that quasistates fit together to form a representation of an interpretation satisfying $\mathcal{O}$, we use the DL-abstraction $\cdot^{\sharp}$ described above, except that now we replace not only nominals but also existential restrictions with fresh concept names. We construct $\mathcal{ML}_{\text{{Diff}}}^{n}$-ontology $\mathcal{O}^{\sharp}$ by replacing CIs $C\sqsubseteq D$ in $\mathcal{O}$ with $C^{\sharp}\sqsubseteq D^{\sharp}$ and extending the result with CIs $\top\sqsubseteq\bigsqcup_{\boldsymbol{T}\in\mathcal{S}_{\mathcal{O}}}\Xi_{\boldsymbol{T}}^{\sharp}$ and $\top\sqsubseteq\exists^{=1}u.\{a\}^{\sharp}$, for all individual names $a$. Then we show that $A$ is $\mathcal{C}$-satisfiable under $\mathcal{O}$ iff ${\boldsymbol{t}}^{\sharp}$ is $\mathcal{C}$-satisfiable under $\mathcal{O}^{\sharp}$, for some type ${\boldsymbol{t}}$ containing $A$. See Appendix C for full details and the case of concept satisfiability (without ontology).
Item (2) is proved by lifting to the modal description logic setting the observation of ? (?) that the difference modality and nominals are mutually interpretable by each other; see the equivalences that show the same expressive power of $\exists^{\neq}$ and $\exists^{=1}$.
Using Theorem 10, one can transfer a large number of (un)decidability results and lower complexity bounds from first-order modal logics with ‘elsewhere’ to modal DLs with non-rigid designators; see, e.g., Theorem 14 in Section 5. Conversely, the results proved below entail new results for first-order modal logics with ‘elsewhere’ and/or counting.
@@ 5 Reasoning in Modal Free Description Logics
The aim of this section is to show the results presented in Table 1 and also discuss a few related logics. For some basic frame classes, it will be convenient to use standard modal logic notation when discussing the satisfiability problem. So, given a propositional modal logic $L$ with $n$ operators and a DL fragment $\mathcal{DL}$, the problem of $L_{\mathcal{DL}}$ concept satisfiability (under global ontology) is the problem of deciding $\mathcal{C}_{L}$-satisfiability of $\mathcal{ML}^{n}_{\mathcal{DL}}$-concepts (under global ontology, respectively), where $\mathcal{C}_{L}$ is the class of all frames validating $L$. We focus on the modal logics $L$ characterised by the following classes $\mathcal{C}_{L}$ of frames, with $n\geq 1$:
@@
$\mathcal{C}_{L}$ is as for $\mathbf{K}^{\ast n}$ and, in addition, $W$ is finite and $R$ is irreflexive (in other words, there are no chains $w_{0}R_{i_{1}}w_{1}\cdots R_{i_{n}}w_{n}$ with $w_{0}=w_{n}$).
@@
For $L\in\{\mathbf{K}^{n},\mathbf{S5}^{n}\}$ with $n\geq 1$, $L_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ concept satisfiability is in NExpTime with both constant and expanding domains.
@@
$\boldsymbol{w}=w_{0}i_{0}w_{1}\cdots i_{m-1}w_{m},$
where $1\leq i_{j}\leq n$, $i_{j}\not=i_{j+1}$, and each $R_{i}$ is the smallest equivalence relation containing all pairs of the form $(\boldsymbol{w},\boldsymbol{w}iw)\in W\times W$. One can assume that $m$ is smaller than the modal depth of $C$. Next, one can substitute the first-order domain by quasistates (as introduced in Section 4) and work with quasimodels $\mathfrak{Q}=(\mathfrak{F},\boldsymbol{q},\mathfrak{R})$, in which $\boldsymbol{q}$ associates a quasistate with any world, and a set of runs $\mathfrak{R}$ represents first-order domain elements as functions mapping every world $w$ to a type in $\boldsymbol{q}(w)$; note that runs were implicit in the proof of Theorem 10 and could be defined as the domain elements in the interpretations for $\mathcal{ML}_{\text{{Diff}}}^{n}$. If ${\boldsymbol{t}}\in\boldsymbol{q}(w)$ contains a nominal, there is only one $r\in\mathfrak{R}$ with $r(w)={\boldsymbol{t}}$; this condition was expressed by the CIs of the form $\top\sqsubseteq\exists^{=1}u.\{a\}^{\sharp}$. Now, one can apply selective filtration and some surgery to such a quasimodel in order to obtain a quasimodel with at most exponential outdegree (and so of at most exponential size), from which one can then extract a model of exponential size.
@@
The proof is again based on appropriate quasimodels, which are now based on expanding domain models using finite frames $\mathfrak{F}=(W,R_{1},\ldots,R_{n},R)$ such that the transitive closure $R$ of $R_{1}\cup\cdots\cup R_{n}$ contains no cycles. Unfolding shows that we may assume that $(W,R)$ is a forest. We show decidability by proving a recursive bound on the size of these models. Let $\mathbb{N}_{\infty}=\mathbb{N}\cup\{\infty\}$, where we assume $m\leq\infty$ and $m+\infty=\infty$, for all $m\in\mathbb{N}_{\infty}$. We fix an ordering ${\boldsymbol{t}}_{1},\ldots,{\boldsymbol{t}}_{k}$ of the types and represent a quasistate as a vector $(x_{1},\ldots,x_{k})\in\mathbb{N}_{\infty}^{k}$, where $x_{i}$ is the number of domain elements that satisfy type ${\boldsymbol{t}}_{i}$ in a world (equivalently, the number of runs through ${\boldsymbol{t}}_{i}$). Let $|\boldsymbol{x}|=x_{1}+\cdots+x_{k}$ for $\boldsymbol{x}=(x_{1},\ldots,x_{k})\in\mathbb{N}_{\infty}^{k}$. Observe that expanding domains correspond to the condition that $wR_{i}v$ implies $|\boldsymbol{x}|\leq|\boldsymbol{y}|$, where $\boldsymbol{q}(w),\boldsymbol{q}(v)$ are represented by $\boldsymbol{x},\boldsymbol{y}\in\mathbb{N}_{\infty}^{k}$, respectively. To obtain a recursive bound on the size of a finite model satisfying a concept we then apply Dickson’s Lemma to the quasistates. Define the product ordering $\leq$ on $\mathbb{N}_{\infty}^{k}$ by setting $\boldsymbol{x}\leq\boldsymbol{y}$ if $x_{i}\leq y_{i}$ for all $1\leq i\leq k$, where $\boldsymbol{x}=(x_{1},\ldots,x_{k})$ and $\boldsymbol{y}=(y_{1},\ldots,y_{k})$. A pair $\boldsymbol{x},\boldsymbol{y}$ with $\boldsymbol{x}\leq\boldsymbol{y}$ is called an increasing pair. Dickson’s Lemma states every infinite sequence $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots\in\mathbb{N}_{\infty}^{k}$ contains an increasing pair $\boldsymbol{x}_{i_{1}},\boldsymbol{x}_{i_{2}}$ with $i_{1}<i_{2}$. In fact, assuming $|\boldsymbol{x}_{i}|\leq|\boldsymbol{x}_{i+1}|$ for all $i\geq 0$ and given recursive bounds on $|\boldsymbol{x}_{1}|$ and $|\boldsymbol{x}_{i+1}|-|\boldsymbol{x}_{i}|$, one can compute a recursive bound on the length of the longest sequence without any increasing pair (?). Now, the proof of a recursive bound on the size of a satisfying model consists in manipulating the quasimodel so that the outdegree of the forest is recursively bounded and Dickson’s Lemma becomes applicable. The expanding domain assumption is crucial for this.
@@
Concept satisfiability for $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{LTL}_{\smash{\mathcal{ALCO}_{\!u}}}$ are polytime-reducible to concept satisfiability for $\mathbf{K}\!\boldsymbol{f}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$ and $\mathbf{K}^{\ast n}_{\smash{\mathcal{ALCO}_{\!u}}}$, respectively, with and without ontology and with both constant and expanding domains.
The proof of this reduction from logics of linear frames with transitive closure to logics of branching frames with transitive closure is not trivial but can be done by adapting the reduction given in the proof of Theorem 6.24 in (?) for product modal logics.
Finally, undecidability of concept satisfiability under global ontology for $\mathbf{K}^{n}$ with $n\geq 1$ follows from undecidability of $\mathcal{ML}_{\text{{Diff}}}^{n}$ (?) using the reduction of Theorem 10. The result for $\mathbf{S5}^{n}$ with $n\geq 2$ can be obtained in a similar way, see, e.g., (?).
Theorem 14.
Concept satisfiability under global ontology is undecidable with constant domains for $\mathbf{S5}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$, with $n\geq 2$, and for $\mathbf{K}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$ with $n\geq 1$.
For many important modal logics the decidability status of modal DLs with non-rigid designators remains open. Most prominently, for the modal logics of (reflexive) transitive frames ${\bf K4}$ (and ${\bf S4}$, respectively), decidability of concept satisfiability with or without ontologies and with expanding or constant domains is open. As a “finitary” approximation and a first step to understand ${\bf K4}$ and ${\bf S4}$, we prove, as a specialisation of the proof of Theorem 12, decidability of concept satisfiability for the Gödel-Löb logic $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ ($\mathbf{GL}$ is the logic of all transitive and Noetherian111$(W,R)$ is Noetherian if there is no infinite chain $w_{0}Rw_{1}R\cdots$ with $w_{i}\not=w_{i+1}$. frames (?)) and Grzegorczyk’s logic $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ ($\mathbf{Grz}$ is the logic of all reflexive and transitive Noetherian frames (?)) in expanding domain models with and without ontologies. Alternatively, the decidability of concept satisfiability for $\mathbf{GL}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ and $\mathbf{Grz}_{\smash{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ can be proved by a double-exponential-time reduction (similar to Theorem 10) to satisfiability in expanding domain products of transitive Noetherian frames, which is known to be decidable (?).
@@ 6 Reasoning in Temporal Free Description Logics
A flow of time $\mathfrak{F}$ is a pair $(T,<)$, where $T$ is either the set $\mathbb{N}$ of non-negative integers or a subset of $\mathbb{N}$ of the form $[0,n]$, for $n\in\mathbb{N}$, and $<$ is the strict linear order on $T$. Elements of $T$ are called instants (rather than worlds). A flow of time $(T,<)$ naturally gives rise to a frame $(T,<,\textit{succ})$ for $\smash{\mathcal{TL}_{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$, where succ is the successor relation: $\textit{succ}=\{(t,t+1)\mid t,t+1\in T\}$. So, we will often say that an interpretation $\mathfrak{M}$ is based on a flow $\mathfrak{F}$ if its frame is induced by $\mathfrak{F}$. We denote it (with an abuse of notation) by $\mathfrak{M}=(\Delta,(\mathcal{I}_{t})_{t\in T})$. If $\mathfrak{M}$ is based on $(\mathbb{N},<)$, then we call it an $\smash{\mathbf{LTL}_{\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ interpretation; if it is based on $([0,n],<)$, for some $n\in\mathbb{N}$, it is called an $\smash{\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\!\mathcal{ALCO}_{\!u}^{\smash{\iota}}}}$ interpretation.
@@
Note that $\Diamond$ is interpreted by $<$ and so does not include the current instant.
@@
$\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}_{\!u}}}$, $\mathbf{LTL}_{\smash{\mathcal{ALCO}_{\!u}}}$ $\Sigma_{1}^{1}$-complete [T 15] undecidable [T 16.1] $\Sigma_{1}^{1}$-complete [T 15] undecidable [T 16.1]
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ALCO}_{\!u}}}$, $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}_{\!u}}}$ undecidable [T 15]
@@
 [T 16.2]
$\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}}}$, $\mathbf{LTL}_{\smash{\mathcal{ALCO}}}$ $\Sigma_{1}^{1}$-complete [T 15] undecidable [T 16.1]
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ALCO}}}$, $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}_{\smash{\mathcal{ALCO}}}$ undecidable [T 15]
@@
 [T 16.2]
$\mathbf{LTL}^{\textstyle\circ}_{\smash{\mathcal{ALCO}_{\!u}}}$ / $\mathbf{LTL}^{\textstyle\circ}_{\smash{\mathcal{ALCO}}}$ Exp-c. / in Exp [T 18] Exp-c. / in Exp [T 18] undecidable [T 19] ?
$\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\textstyle\circ}_{\smash{\mathcal{ALCO}_{\!u}}}$ / $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\textstyle\circ}_{\smash{\mathcal{ALCO}}}$ Exp-c. / in Exp [T 18] Exp-c. / in Exp [T 18] undecidable [T 19] decidable
@@
We begin the study of the satisfiability problems for temporal DLs based on $\smash{\mathcal{ALCO}}$ by showing that concept satisfiability in constant domains is undecidable or even $\Sigma_{1}^{1}$-complete (highly undecidable in the analytical hierarchy) over the infinite flow of time $(\mathbb{N},<)$. This is very different from the classical case with RDA, which was shown to be decidable in the absence of definite descriptions (?, Theorem 14.12).
@@
The lower bounds follow, using Theorem 10 (2), from the respective undecidability results for the first-order one-variable temporal logic with counting (?). Observe that, for $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}}}$, we apply the spy-point universal role elimination of Lemma 6. Moreover, Proposition 3 also applies to $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ALCO}}}$ under global ontology, and so the results hold in both total and partial interpretations.
They could also be proven more directly by encoding the $\Sigma_{1}^{1}$-complete recurrence and undecidable reachability problems for the Minsky counter machines (??). Intuitively, the value of a counter of the Minsky machine can be represented as the cardinality of a certain concept. Then non-rigid nominals (or indeed the counting quantifier) can be used to ensure that the value of the counter is incremented/decremented (depending on the command) by the transition: for instance, a CI of the form
@@
(1) CI entailment with constant domains is $\Pi_{1}^{1}$-complete for $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ELO}}}$ and undecidable for $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ELO}}}$.
(2) CI entailment with expanding domains is undecidable for $\mathbf{LTL}^{\Diamond}_{\smash{\mathcal{ELO}}}$.
(3) CI entailment with expanding domains is decidable but Ackermann-hard for $\mathbf{LTL}\!{\text{$\boldsymbol{f}$}}^{\Diamond}_{\smash{\mathcal{ELO}}}$.